Stacy is selling candy bars. She sold 28 on the 4th day, she sold 34 on the 5th day and she sold 40 on the 6th day. Use an algebraic expression to determine how many candy bars she will sell on the 15th day.
Answer by checkley77(12844)

N=A+(N-1)D
N=THE Nth NUMBER.
A=THE FIRST NUMBER
D=THE DIFFERENCE
4th=28
5th=34
6th=40
The difference=+6.
The first number (A)=28-3*6)=28-18=10.
Thus the 15th term is:
15th=10+(15-1)*6)
15th=10+14*6
15th=10+84=94 ans.
Proof:
10,16,22,28,34,40,46,52,58,64,70,76,82,88,94